# Empirical analysis of Binding Precedent efficiency in the Brazilian Supreme Court via Similar Case Retrieval

### v2

@@ 4.7 Final legal considerations
However, the main justifications for the introduction of the BP in the Brazilian constitutional system are predicated on empirical claims. After all, the instrument results from an effort to standardize constitutional jurisprudence and contain the demands that reach the STF, given the high volume of litigation the court must handle every year. This effort, in fact, not only led to the introduction of the BP but also to other instruments aimed at reducing the number of cases judged by the STF. Despite this fact, there are few empirical studies that seek to verify whether the BPs produce the expected effects — particularly studies that aim to ascertain whether there is indeed a standardization of jurisprudence and a reduction in litigation concerning the subject of the BP. The reasons for this are partly methodological, as discussed earlier in this text.
@@
The first and most frequent explanatory hypothesis for the increase in litigation following the creation of an BP is what we refer to as the creation of a New avenue to STF or procedural issues. This hypothesis is relevant for explaining almost all the cases studied [^41] It is less present only in BP 17, but was explicitly identified as a relevant cause of litigation in BPs 11, 14, 26, and 37.. This arises from the fact that the judicial system involves actors who act strategically. It is often a mistake to assume, when seeking a social outcome through a legal change, that the recipients of that change will continue to behave the same way they did before its introduction. Generally, it is necessary to understand how a change in the law will alter the incentives of the actors involved. The claim that a judicial decision is not in accordance with an BP allows the case to reach the STF, creating an interest among litigants to address this type of issue precisely to gain access to the court. In other words, by establishing an BP on a particular subject, the STF increases the constitutional relevance of the issue, signaling to litigants that it will hear cases involving violations of that BP.
Therefore, the establishment of an BP always creates incentives for parties to invoke the issue of the BP to access the court, even if their cases’ connection to the subject of the BP is weak or indirect. The same process, in fact, can be observed with the creation of legal rules through the usual legislative process: it is a well-known legal phenomenon that the multiplication of normative acts establishing specific norms tends to increase the complexity of the existing legal system, generating more legal uncertainty and litigation, instead of providing clarity to the law applicable to each specific case. The language of the BP itself can introduce new complexities if the way the BP addresses its subject contains ambiguities or vague terms that give rise to new judicial discussions.
@@
The semantic openness of normative statements is also what allows groups of litigants to resist the application of an BP or for divergent interpretations to arise in specific lower courts, creating regional specificities in how the BP is addressed. As a rule, these points of resistance need to mobilize some ambiguity or legal gap left by the BP. The hypothesis we called Resistance by a group of litigants or regional specificity proved relevant in two cases: BP 26 and BP 37. In these cases, we see that the creation of an BP will not always have the effect of pacifying social relations and reducing conflicts; it can also incite more litigation and opposition from those affected, especially when the content of the BP does not seem to adequately address a particular case and the agents seek to differentiate it from the original scope of application of the BP (the well-known distinguishing strategy).
@@
All five hypotheses listed in this study suggest that the establishment of an BP can, in certain cases — or possibly in most cases — result in an increase in the number of litigations, raising new controversies and legal issues that demand judicial solutions. This problem has already been discussed theoretically by legal doctrine, but it has not yet been measured, despite being based on claims that are, as stated, essentially empirical. We hope that the methodology proposed here will provide important contributions to a better understanding of the impacts of BPs and the possible causes of increased litigation after the issuance of an BP.
@@
It is important to note that our entire analysis, which claims to uncover the mechanisms underlying the repetitive use of a binding precedent, is significantly biased by the fact that we perform Similar Case Retrieval only on the documents brought up to trial at the Supreme Court. In particular, our method is vulnerable to a survivorship bias, unaware of themes that are confined to the lower courts, or not taken to court at all (e.g., theses from advocacy groups, reports from various agencies, political events).
This problem can be mitigated in at least three ways. The first, already mentioned above, would be to acquire data from courts below the STF. In the case of STJ (Superior Court of Justice), data is available on the official website [^42] https://www.stj.jus.br/sites/portalp/Processos/Consulta-Processual. This way, the authors of [41] extracted 129,602 cases, enabling an analysis of corporate law precedents. For other courts, however, a search engine is not available, and we are not aware of any public database — the data acquisition must be done by asking each court.
Another important avenue would be a field survey, involving interviews with legal professionals (e.g., lawyers, public defenders, prosecutor, or judges). In this regard, we had the opportunity to speak with members of the Public Defender’s Office and Public Prosecutor’s Office of the State of São Paulo, who kindly presented to us, from their point of view, the important issues at stake in the creation of the BPs (more specifically BP 26). However, the development of a sociological analysis, which would take into account the various political and legal positions with regard to Supreme Court precedents, would require the collection of many other points of view.
Finally, our analysis would benefit from the incorporation of expert annotations in the training set, in order to isolate key elements of the documents. Indeed, our models are trained in such a way that all parts of the text are considered equally; this potentially introduces an important bias, by encouraging the integration of elements that are too specific to certain cases, or by giving equal importance to cases judged solely on procedural grounds. Such an approach is expected to refine our models, which suffer, as our work has shown, from a considerable generalization error. In this regard, and given the considerable number of documents studied, it would be interesting to produce automatic annotations using Large Language Models.
@@ 5 Conclusion
In this article, we have performed a series of mathematical and juridical analyses that allowed us to explain why BPs 11, 14, 17, 26, and 37 (some of the most cited in the STF’s decisions) fail in reducing STF’s workload. First, we have applied and compared different Similar Case Retrieval methods (TF-IDF-based models, LSTM, BERT, and regular expressions) on a database composed of STF’s decisions. By studying time series of similar cases, our methodology enabled us to assess the impact of laws on jurisprudence, and to perform an empirical study of the juridical mechanisms behind these BPs’ inefficiency. We finally described five main hypotheses that explain the large number of cases reaching STF.
From the mathematical point of view, the TF-IDF models used in our analyses performed better than LSTM and BERT. This was observed both on the annotated dataset (Dataset #1), where the scores of the first models are higher, and on the larger dataset (Dataset #2), where the deep learning models face a considerable generalization error. Although we have used a Portuguese-specific BERT model, fine-tuned on a large collection of legal documents, we plan to incorporate further improvements: circumventing the truncation of input text to the first 510/512 tokens (e.g., via pooling [76] or Big Bird [74]), and fine-tuning to our specific dataset. In addition, we intend to investigate whether the incorporation of expert-annotated key parts would lead to better results or new insights.
From a legal perspective, our study uncovers a counter-intuitive effect, that has been overlooked in purely doctrinal works: the very nature of the binding precedent brings with it reasons for increasing repetitive demands. This includes the fact that, when a BP is published, it is declared to be a new topic of relevance that the Supreme Court is committed to dealing with; that it can lead to the formation of diverse groups of litigants; that it introduces new complexities and potentially ambiguities; that it exposes and legitimizes a new discussion topic; and that it is naturally subject, in an indirect way, to other economic, political or social conjectures. In order to refine our study, which is based solely on the analysis of STF decisions, we plan to integrate somehow lower court decisions, as well as conducting a field survey to gather the views of various legal professionals.

### v3

@@ 5.7 Final legal considerations
However, the main justifications for the introduction of the BP in the Brazilian constitutional system are predicated on empirical claims. After all, the instrument results from an effort to standardize constitutional jurisprudence and contain the demands that reach the STF, given the high volume of litigation the court must handle every year. This effort, in fact, not only led to the introduction of the BP but also to other instruments aimed at reducing the number of cases judged by the STF. Despite this fact, few empirical studies seek to verify whether the BPs produce the expected effects — particularly studies that aim to ascertain whether there is indeed a standardization of jurisprudence and a reduction in litigation concerning the BP’s subject. The reasons for this are partly methodological, as discussed earlier in this text.
@@
The first and most frequent explanatory hypothesis for the increase in litigation following the creation of a BP is what we refer to as the creation of a New avenue to STF or procedural issues. This hypothesis is relevant for explaining almost all the cases studied [^55] It is less present only in BP 17, but was explicitly identified as a relevant cause of litigation in BPs 11, 14, 26, and 37.. This arises from the fact that the judicial system involves actors who act strategically. It is often a mistake to assume, when seeking a social outcome through a legal change, that the recipients of that change will continue to behave the same way they did before its introduction. Generally, it is necessary to understand how a change in the law will alter the incentives of the actors involved. The claim that a judicial decision is not in accordance with a BP allows the case to reach the STF, creating an interest among litigants to address this type of issue precisely to gain access to the court. In other words, by establishing a BP on a particular subject, the STF increases the constitutional relevance of the issue, signaling to litigants that it will hear cases involving violations of that BP.
Therefore, the establishment of a BP always creates incentives for parties to invoke the issue of the BP to access the court, even if their cases’ connection to the subject of the BP is weak or indirect. The same process, in fact, can be observed with the creation of legal rules through the usual legislative process: it is a well-known legal phenomenon that the multiplication of normative acts establishing specific norms tends to increase the complexity of the existing legal system, generating more legal uncertainty and litigation, instead of providing clarity to the law applicable to each specific case. The language of the BP itself can introduce new complexities if the way the BP addresses its subject contains ambiguities or vague terms that give rise to new judicial discussions.
@@
The semantic openness of normative statements is also what allows groups of litigants to resist the application of a BP or for divergent interpretations to arise in specific lower courts, creating regional specificities in how the BP is addressed. As a rule, these points of resistance need to mobilize some ambiguity or legal gap left by the BP. The hypothesis we called Resistance by a group of litigants or regional specificity proved relevant in two cases: BP 26 and BP 37. In these cases, we see that the creation of a BP will not always have the effect of pacifying social relations and reducing conflicts; it can also incite more litigation and opposition from those affected, especially when the content of the BP does not seem to adequately address a particular case and the agents seek to differentiate it from the original scope of application of the BP (the well-known distinguishing strategy).
@@
All five hypotheses listed in this study suggest that the establishment of a BP can, in certain cases — or possibly in most cases — result in an increase in the number of litigations, raising new controversies and legal issues that demand judicial solutions. This problem has already been discussed theoretically by legal doctrine, but it has not yet been measured, despite being based on claims that are, as stated, essentially empirical. We hope that the methodology proposed here will provide important contributions to a better understanding of the impacts of BPs and the possible causes of increased litigation after the issuance of a BP.
@@
It is important to note that our entire analysis, which claims to uncover the mechanisms underlying the repetitive use of a binding precedent, is significantly biased by the fact that we perform Case Classification only on the documents brought up to trial at the Supreme Court. In particular, our method is vulnerable to a survivorship bias, unaware of themes that are confined to the lower courts, or not taken to court at all (e.g., theses from advocacy groups, reports from various agencies, political events).
This problem can be mitigated in at least three ways. The first, already mentioned above, would be to acquire data from courts below the STF. In the case of STJ (Superior Court of Justice), data is available on the official website [^56] STJ search engine https://www.stj.jus.br/sites/portalp/Processos/Consulta-Processual. This way, the authors of [42] extracted 129,602 cases, enabling an analysis of corporate law precedents. For other courts, however, a search engine is not available, and we are not aware of any public database — the data acquisition must be done by asking each court.
Another important avenue would be a field survey, involving interviews with legal professionals (e.g., lawyers, public defenders, prosecutors, or judges). In this regard, we had the opportunity to speak with members of the Public Defender’s Office and Public Prosecutor’s Office of the State of São Paulo, who kindly presented to us, from their point of view, the important issues at stake in the creation of the BPs (more specifically BP 26). However, the development of a sociological analysis, which would take into account the various political and legal positions with regard to Supreme Court precedents, would require the collection of many other points of view.
Finally, our analysis would benefit from the incorporation of expert annotations in the training set, in order to isolate key elements of the documents. Indeed, our models are trained in such a way that all parts of the text are considered equally; this potentially introduces an important bias, by encouraging the integration of elements that are too specific to certain cases, or by giving equal importance to cases judged solely on procedural grounds. Such an approach is expected to refine our models. In this regard, and given the considerable number of documents studied, it would be interesting to produce automatic annotations using Large Language Models.
@@ 6 Conclusion
In this article, we have performed a series of mathematical and juridical analyses that allowed us to explain why BPs 11, 14, 17, 26, and 37 (some of the most cited in the STF’s decisions) fail in reducing the Court’s workload. First, we have applied and compared different Case Classification methods (TF-IDF-based models, LSTM, Longformer, and regular expressions) on a database composed of STF’s decisions. By studying time series of similar cases, our methodology enabled us to assess the impact of laws on jurisprudence and to perform an empirical study of the juridical mechanisms behind these BPs’ inefficiency. We finally described five main hypotheses that explain the large number of cases reaching STF.
From the mathematical point of view, the TF-IDF models used in our analyses performed slightly better than LSTM and Longformer, when compared through the $F_{1}$ score calculated from the labeled dataset (Dataset #1). Nevertheless, this is counterbalanced by the fact that, in the larger dataset (Dataset #2), the deep learning models were able to detect certain important legal events that TF-IDF missed. To enable a more refined analysis of the predictions, it would be interesting to consider models that not only detect cases falling under the BP’s field of application but also the role it could play in the case at hand. As we have seen throughout the article, binding precedents can be used for a wide variety of purposes, whether for supporting or countering an argument, providing context, or introducing a legal doctrine, but they are also used to judge cases based on their merits or procedural grounds, an important piece of information for characterizing the applications of the BP. In addition, we intend to investigate whether the incorporation of expert-annotated key parts would lead to better results or new insights.
From a legal perspective, our study uncovers a counter-intuitive effect that has been overlooked in purely doctrinal works: the very nature of the binding precedent brings with it reasons for increasing repetitive demands. This includes the fact that, when a BP is published, it is declared to be a new topic of relevance that the Supreme Court is committed to dealing with; that it can lead to the formation of diverse groups of litigants; that it introduces new complexities and potentially ambiguities; that it exposes and legitimizes a new discussion topic; and that it is naturally subject, in an indirect way, to other economic, political or social conjectures. To refine our study, which is based solely on the analysis of STF decisions, we plan to integrate somehow lower court decisions, as well as conduct a field survey to gather the views of various legal professionals.